If 15 is subtracted from twice a number, the difference is five times the number
Answer by checkley77(12844)

2X-15=5X
2X-5X=15
-3X=15
X=-15/3
X=-5 ans.
Proof:
2*-5-15=5*-5
-10-15=-25
-25=-25